OpenStudy (anonymous):

If I invest $2,000 on 1/1/01 and it returns $10,000 on 1/1/06, what is the rate of return?

OpenStudy (anonymous):

You could use the future value formula: \[FV =P(1+r)^{Y}\] Where FV is the future value of your money ($10,000), P is the principle, or initial amount of money ($2,000), r is the annual rate of return (which we're trying to solve for), and Y is the number of years the money is let to grow (2006-2001 = 5 years). So when you plug in these numbers the equation looks like this: \[10,000=2,000(1+r)^{5}\] And then solve for r: \[[10,000=2,000(1+r)^{5}]/2,000\rightarrow5=(1+r)^{5}\] \[\sqrt[5]{5}=\sqrt[5]{(1+r)^{5}}\rightarrow1.379723=1+r\] \[1.37973-1=1+r-1\rightarrow r=0.37973=37.973percent\] So your annual rate of return is 37.973 %.

OpenStudy (anonymous):

Jpomer35 - that makes sense to me. Thanks so much for your help!